OpenStudy (anonymous):

Solve for x: the quantity of x plus 4 over 3 = 2. x = -2 x = 2 x = 2 over 3 x = negative 10 over 3

OpenStudy (shamil98):

multiply 3 to both sides to get rid of the fraction, then subtract 4 from both sides to get the value of x.

OpenStudy (anonymous):

the value of x is ?

OpenStudy (shamil98):

\[\frac{ x+4 }{ 3 } = 2\] \[x + 4 = 6\] x = ?

OpenStudy (shamil98):

positive 2

OpenStudy (anonymous):

thank you